# Discovering Message Passing Hierarchies for Mesh-Based Physics Simulation

### v1

@@ 3.4 Inter-Level Feature Propagation with AMP
Prior works employed non-parametric aggregation in inter-level propagation, convolving features based on the normalized node degree. It simplifies intricate relationships between nodes and neglects the directional aspects of information flow. In comparison, the inter-level aggregation weights in DHMP are data-specific and time-varying. Notably, the importance weight $\alpha^{l}_{ij}$ in the proposed AMP layer inherently captures the significance of node $v_{j}$’s features to node $v_{i}$ at the graph level $l$. Consequently, it can be directly reused for the REDUCE and EXPAND operations in the downsampling and upsampling processes. We provide details of these operations as follows:
  • REDUCE: Let $v_{i}$ be the node at the coarser graph level. The downsampling process aggregates the information of the current neighbors $\mathcal{N}_{i}$ by reusing the weight $\alpha^{l}_{ij}$: $\mathbf{v}_{i}^{l+1}\leftarrow\sum_{j\in\mathcal{N}_{i}}\alpha^{l}_{ij}\mathbf{v}^{l}_{j}$.
  • EXPAND: We first unpool the node features from the next-level coarser graph $\mathcal{G}_{l+1}$. To achieve this, we record the nodes selected during the downsampling process and use this information to place nodes back in their original positions in the graph. Next, we re-use the importance weight $\alpha^{l}_{ij}$ to assign features in the coarser graph to nodes in the finer graph, i.e., $\mathbf{v}^{l}_{i}\leftarrow\sum_{j\in\mathcal{N}_{i}}\mathbf{v}^{l+1}_{j}\alpha^{l}_{ij}$.
  • FeatureMixing: Following the EXPAND operation, DHMP conducts an additional message passing step based on $\mathbf{v}^{l}_{i}$. It then integrates the resulting features with the intra-level message passing outcomes in $\mathcal{G}_{l}$ (before downsampling) through a skip connection.
@@ 3.5 Implementation Details
We train DHMP using the one-step supervision that measures the $L_{2}$ loss between the ground truth and the next-step predictions. We include detailed descriptions of the physical quantities represented by input node and edge features in Appendix B. We implement the encoder, decoder, node update function $\phi^{v}$, and edge update function $\phi^{e}$ using two-layer MLPs with ReLU activation and a hidden size of $128$. Likewise, the network component for generating importance weights, $\phi^{w}$, in AMP is implemented using a two-layer MLP. We apply layer normalization to the MLP outputs, except for those of the decoder and the importance weight network. We discuss the specific number of graph levels $L$ for downsampling in Appendix A. In the Gumbel-Softmax operator for differentiable node selection, we use temperature annealing to decrease the temperature from $5$ to $0.1$ with a decay factor of $\gamma=0.999$, which aims to encourage the exploration of hierarchies while gradually refining node selection to ensure training stability.
@@ 4 Experiments
Table 2: Quantitative comparison of the one-step and long-term prediction errors. We report the mean results over $3$ random seeds, with corresponding standard deviations detailed in Appendix F. Promotion denotes the improvement over the second-best model.
Model RMSE-1 (×$\times 10^{-2}$) RMSE-All (×$\times 10^{-2}$)
Cylinder Airfoil Flag Plate Cylinder Airfoil Flag Plate
MGN (2021) 0.4046 77.38 0.4890 0.0579 59.78 2816 124.5 3.982
BSMS-GNN (2023) 0.2263 71.69 0.5080 0.0632 16.98 2493 168.1 1.811
Lino et al. (2022) 3.9352 85.66 0.9993 0.0291 27.60 2080 118.2 2.090
HCMT (2024) 0.9190 48.62 0.4013 0.0295 23.59 3238 90.32 2.468
DHMP 0.1568 41.41 0.3049 0.0282 6.571 2002 76.16 1.296
Promotion 30.7% 14.8% 24.0% 3.10% 61.3% 3.75% 15.7% 28.5%
@@ 4.1 Experimental Setup
Datasets.
We evaluate our approach on five mesh-based physics simulation benchmarks established in previous literature (Pfaff et al., 2021; Cao et al., 2023; Wu et al., 2023; Narain et al., 2012).
@@
  • Flag: Simulation of flag dynamics in the wind based on Lagrangian meshes with fixed topology.
@@
  • FoldingPaper: Deformation of paper sheets on Lagrangian meshes, with varying forces at the four corners and evolving mesh graph.
For details regarding the datasets, including descriptions of the input physical quantities, please refer to Appendix A. Additional information concerning our implementation can be found in Appendix B.
Compared models.
We primarily compare DHMP with the following methods:
  • MGN (Pfaff et al., 2021), which performs $15$ times of message passing at the original graph level.
@@
  • Lino et al. (2022), which also trains MPNNs on manually-set multi-scale mesh graphs.
@@
All models are trained using the Adam optimizer with an exponential learning rate decay from $10^{-4}$ to $10^{-6}$. Detailed descriptions and hyperparameters of the models are provided in the Appendix C.
@@ 4.2 Main Results
Table 2 presents the root mean squared error (RMSE) of one-step prediction (RMSE-1) and long-term rollouts for 100100100100600600600600 future time steps (RMSE-all). DHMP consistently outperforms the compared models across all benchmarks. This demonstrates the effectiveness of building context-aware, temporally evolving hierarchies with learnable, directionally non-uniform feature propagation both within and across graph levels. Figure 2 presents long-term predictions on CylinderFlow, based solely on the system’s initial conditions at the first step. As we can see, DHMP captures the complex, time-varying fluid flow around the cylinder obstacle more successfully, with its predictions closely matching the ground truth evolution. More results are shown in Appendix I.
@@
In Figure 3, we visualize the dynamic hierarchies constructed by DHMP at different time steps, where coarser-level nodes tend to concentrate in regions highlighted by the temporal differences in the true data. We have two observations here: First, the constructed hierarchy evolves as the data context changes. Second, the temporally evolving graph structures align with the high-intensity regions, either in the velocity fields (top) or in their temporal variations (bottom). These findings highlight the effectiveness of our approach in capturing significant patterns within the dynamic system.
Figure 2: Prediction showcases over $400$ future steps on CylinderFlow and the final error maps.
Figure 3: Top: the velocity field from the true data. Bottom: the temporal difference of the velocity fields between adjacent time steps alongside the constructed coarser-level mesh graph ($\mathcal{G}_{l=4}$). The highlighted areas demonstrate a notable experimental phenomenon: the mesh dynamically evolves with the data context, and aligns with the critical areas of change in the data.
Table 3: Errors of 2D paper simulation (×$\times 10^{-2}$) on time-varying meshes.
Model RMSE-1 RMSE-All
MGN 0.0618 24.08
DHMP 0.0544 7.41
@@
We evaluate DHMP in a more challenging setting with time-varying meshes for paper folding simulation, generated using the ARCSim solver (Narain et al., 2012; Wu et al., 2023). Notably, methods such as BSMS-GNN and HCMT require pre-computed hierarchies as part of their data preprocessing, which poses a significant limitation in scenarios with continuously changing mesh topologies. We assess the models using ground-truth remeshing nodes provided by the ARCSim Adaptive Remeshing component, following the setup from Pfaff et al. (2021). As shown in Table 3, DHMP achieves superior short-term and long-term accuracy compared to MGN, indicating that the dynamic graph hierarchies in our approach can better fit physical systems with significant geometric variations, as represented by the time-varying input mesh structures.
@@
Due to the stochasticity of Gumbel-Softmax sampling in DiffSELECT, we evaluate the stability of trained DHMP by conducting three independent runs on the test set. The mean and standard deviations of the prediction errors reveal minimal discrepancies across different runs, as shown in Table 10 in Appendix E. These findings demonstrate that once trained, DHMP generates consistent graph hierarchies based on the same inputs.
Computation efficiency.
The computation efficiency is evaluated in Appendix G. It shows that DHMP has the lowest training cost and lower inference time compared to attention-based model.
@@ 4.3 Ablation Studies
DHMP has three contributions: (i) dynamic hierarchy, (ii) anisotropic intra-level propagation, (iii) learnable inter-level propagation. To investigate the effectiveness of each component, we implement various variants of DHMP, including Static-Anisotropic-Unlearnable (M1), Static-Anisotropic-Learnable (M2); Dynamic-Anisotropic-Unlearnable (M3), and compare them against BSMS-GNN. The baseline model uses static hierarchies, isotropic intra-level summation, and unlearnable inter-level propagation. Ablation study results compared to baseline BSMS-GNN are presented in Figure 4.
@@
Figure 4: The analyses of dynamic hierarchies, anisotropic intra-level propagation, and learnable inter-level feature propagation. The red dashed lines represent results from BSMS-GNN (Cao et al., 2023). Lower values indicate better performance.
(a) Error distribution for BSMS-GNN and DHMP
(b) Ratios of challenging nodes
Figure 5: (a) Error maps, where nodes with the top $10\%$ of errors in each model’s predictions are marked in yellow and referred to as “challenging nodes”. (b) DHMP retains more challenging nodes in coarser graph hierarchies to capture multi-scale dependencies more effectively.
Effectiveness of dynamic hierarchies.
From Figure 4, by comparing DHMP vs. M2 and M3 vs. M1, we observe the advantages of learning dynamic hierarchical graph structures. These results highlight the significance of adaptively modeling interactions in context-dependent graphs. To better understand how DHMP constructs dynamic hierarchies, we visualize the distribution of nodes with the top $10\%$ prediction errors in Figure 5(a). Accordingly in Figure 5(b), we observe that DHMP retains a higher proportion of “challenging” nodes in the coarser message passing levels, enabling our model to capture multi-scale dependencies more effectively, especially in areas where finer message passing levels struggle. In contrast, the predefined static hierarchies in the BSMS baseline are data-independent and may inevitably overlook modeling long-range relations surrounding these pivotal nodes, even though they typically present higher errors than those in DHMP.
Effectiveness of anisotropic message passing.
Figure 4 further illustrates the importance of enhancing the direction-specific contributions during both intra-level and inter-level updates. First, incorporating AMP into the static hierarchy results in performance improvements, as shown by the comparison between M1 and BSMS-GNN. Additionally, the significance of transmitting directed inter-level information is highlighted by comparing DHMP vs. M3 and M1 vs. M2. In Figure 6, we visualize the variance of predicted anisotropic edge weights and compare it with areas where physical quantities present substantial variations over time. The results reveal a strong correlation between the anisotropic learning mechanism and the rapidly changing dynamics of the physical system.
Figure 6: Visualizations of the variance of the generated anisotropic weights calculated on adjacent edges (top) and the corresponding variance of physical quantities computed over time (bottom). The strong correlation between them reveals the AMP’s ability to perceive significant patterns in data.
4.4 Generalization Analyses
Generalization to out-of-distribution mesh resolutions.
Almost none of the existing machine learning models for mesh-based physics simulation are resolution-free. They may fail when evaluated on unseen mesh resolutions. We assess the generalization performance of DHMP by training it on low-resolution meshes and testing it on high-resolution meshes. The average number of nodes in the test data is twice that of the training data, and the number of edges is three times greater. As shown in Table 4, DHMP demonstrates improved zero-shot generalization ability to more refined mesh structures. This improvement is primarily attributed to our model’s capability to generate hierarchical graphs adaptively. However, it is important to note that this result does not imply that our method has fully explored generalization across arbitrary resolutions—achieving true resolution-free modeling requires a more refined model design. Nevertheless, this holds significant value in practical applications and has the potential to greatly reduce the time overhead of numerical simulation processes for preparing the large-scale mesh data required for model training.
Table 4: Results on out-of-distribution (OOD) mesh resolutions.
Model RMSE-1 (×$\times 10^{-2}$) RMSE-All (×$\times 10^{-2}$)
@@ 4.4 Generalization Analyses
BSMS-GNN 0.9177 202.3 0.6486 0.0474 33.87 6179 148.2 1.904
DHMP 0.4855 126.7 0.5536 0.0368 47.72 5759 120.9 2.553
Table 5: Generalization results across various scales of input velocities, presented by the variance and norm of data in training/test splits. Increase denotes the relative increase of the test data compared to the training data.
Cylinder Airfoil
Split Var Norm Var Norm
@@
Model RMSE-1 RMSE-All RMSE-1 RMSE-All
@@
DHMP 2.14×$\times 10^{-3}$ 0.091 0.665 22.57
Generalization to physics variations.
We evaluate DHMP under strong distribution shifts in the input physical quantities. Table 5 presents data statistics and the RMSE results on the CylinderFlow and Airfoil datasets. DHMP achieves lower RMSEs than BSMS-GNN in both short-term and long-term simulations, which can be largely attributed to the proposed AMP layer. When the fluid dynamics in the test set become more complex—characterized by increased variance in the velocity field over time—the dynamics patterns propagate more rapidly in space. The AMP layer can more effectively capture directed long-range node interactions.

### v2

@@ 3.3 Inter-Level Feature Propagation with AMP
Prior work employed non-parametric aggregation in inter-level propagation, convolving features based on the normalized node degree. It simplifies intricate relationships between nodes and neglects the directional aspects of information flow. To address this, EvoMesh is designed to learns inter-level aggregation weights that are both data-specific and time-varying. Specifically, the importance weight $\alpha^{l}_{ij}$ computed by the AMP layer inherently captures the relevance of node $v_{j}$’s features to node $v_{i}$ at the graph level $l$. These weights can be directly reused for the REDUCE and EXPAND operations in the downsampling and upsampling processes. We provide details of these operations as follows:
  • REDUCE: Let $v_{i}$ be the node at the coarser graph level. The downsampling process aggregates the information of the current neighbors $\mathcal{N}_{i}$ by reusing the weight $\alpha^{l}_{ij}$: $\mathbf{v}_{i}^{l+1}\leftarrow\texttt{REDUCE}(\{\mathbf{v}_{j}^{l},\alpha^{l}_{ij}\}_{j\in\mathcal{N}_{i}}):=\sum_{j\in\mathcal{N}_{i}}\alpha^{l}_{ij}\,\mathbf{v}_{j}^{l}.$
  • EXPAND: We first unpool the node features from the coarser graph $\mathcal{G}_{l+1}$ back to the finer level $\mathcal{G}_{l}$. To achieve this, we record the nodes selected during the downsampling process and use this information to place the nodes back in their original positions in the graph. Then, we reuse the previously computed importance weights $\alpha_{ij}^{l}$ to assign weighted features from $\mathcal{G}_{l+1}$ back to $\mathcal{G}_{l}$. The EXPAND operation is formally defined as: $\tilde{\mathbf{v}}_{i}^{l}\leftarrow\texttt{EXPAND}(\{\mathbf{v}_{j}^{l+1},\alpha^{l}_{ij}\}_{j\in\mathcal{N}_{i}}):=\sum_{j\in\mathcal{N}_{i}}\mathbf{v}_{j}^{l+1}\alpha^{l}_{ij}.$
  • FeatureMixing: While the EXPAND operation upsamples coarser-level features of $G_{l+1}$ to match the resolution of the current level $G_{l}$, naïvely upsampled features may suffer from artifacts or misalignment. To mitigate this, we introduce FeatureMixing to refine and fuse coarse-level features with intra-level information at the current resolution. Specifically, we apply an additional anisotropic message passing step to the upsampled features $\tilde{\mathbf{v}}^{l}_{i}$ and then integrate these features with the original intra-level features $\mathbf{v}_{l}$ in $\mathcal{G}_{l}$ (prior to downsampling) using a skip connection: $\bar{\mathbf{v}}_{i}^{l}\leftarrow\texttt{FeatureMixing}(\tilde{\mathbf{v}}_{i}^{l},\mathbf{v}_{i}^{l},\{\mathbf{e}_{ij}^{l}\}_{j\in\mathcal{N}_{i}}):=\mathbf{v}_{i}^{l}+\texttt{AMP}(\tilde{\mathbf{v}}_{i}^{l},\{\tilde{\mathbf{v}}_{j}^{l}\}_{j\in\mathcal{N}_{i}},\{\mathbf{e}^{l}_{ij}\}_{j\in\mathcal{N}_{i}}).$
@@ 3.4 Implementation Details
We train EvoMesh using the one-step supervision that measures the $L_{2}$ loss between the ground truth and the next-step predictions. We include detailed descriptions of the implementation of encoder, decoder, node update function and edge update function in Appendix B.
@@ 4 Experiments
In this section, we present the key results of the proposed method. Additional analyses can be found in Appendix C.
Figure 2: Prediction showcases over 400 future steps on CylinderFlow. From the displayed error maps, it is evident that EvoMesh effectively captures long-term dynamics, providing predictions that closely align with the ground truth.
@@ 4.1 Experimental Setup
We evaluate EvoMesh on five mesh-based benchmarks from previous work (Pfaff et al., 2021; Cao et al., 2023; Wu et al., 2023; Narain et al., 2012). For detailed descriptions, including the input physical quantities, please refer to Appendix A.
@@
  • FlyingFlag: Simulation of flag dynamics in the wind based on Lagrangian meshes with fixed topology.
@@
  • FoldingPaper: Deformation of paper sheets with evolving meshes driven by varying forces at the four corners.
We mainly compare EvoMesh with the following methods:
  • MGN (Pfaff et al., 2021), which performs multiple times of message passing on the original graph.
  • Lino et al. (Lino et al., 2022), which also trains MPNNs on manually-set multi-scale mesh graphs.
@@
  • Eagle (Janny et al., 2023), which constructs a two-scale hierarchy using precomputed geometric clustering and performs message passing at both levels.
@@
All models are trained using the Adam optimizer for $1M$ steps, with an exponential learning rate decay from $10^{-4}$ to $10^{-6}$ over the first $500K$ steps. We provide further details on the architecture and hyperparameters of the compared models in Appendix D.
@@ 4.2 Main Results
Table 2 presents the root mean squared error (RMSE) of one-step prediction (RMSE-1) and long-term rollouts for 100100100100600600600600 future time steps (RMSE-all). EvoMesh consistently outperforms the compared models across all benchmarks. This demonstrates the effectiveness of building context-aware, time-evolving hierarchies with learnable, directionally non-uniform feature propagation both within and across graph levels. Figure 2 presents long-term predictions on CylinderFlow, based solely on the system’s initial conditions at the first step. As we can see, EvoMesh captures the complex, time-varying fluid flow around the cylinder obstacle more successfully, with its predictions closely matching the ground truth evolution. More results are shown in Appendix C.9.
Figure 3: A demonstration of how the learned hierarchies adapt to evolving physical dynamics. Top: the velocity field from the true data. Bottom: the temporal difference of the velocity fields between adjacent time steps alongside the constructed coarser-level mesh graph ($\mathcal{G}_{l=4}$). The highlighted areas demonstrate a notable experimental phenomenon: the mesh dynamically evolves with the data context and aligns with the critical areas of change in the data.
@@
In Figure 3, we visualize the time-evolving hierarchies constructed by EvoMesh at different time steps, where coarser-level nodes tend to concentrate in regions highlighted by the temporal differences in the true data. We have two observations here: First, the constructed hierarchy evolves as the data context changes. Second, the time-evolving graph structures align with the high-intensity regions, either in the velocity fields (top) or in their temporal variations (bottom). These results highlight the effectiveness of our approach in capturing significant dynamic patterns.
Table 3: Simulation results of 2D paper folding with time-varying input meshes. We here compare EvoMesh with MGN, as BSMS and HCMT require pre-computed hierarchies, which are unsuitable for scenarios involving continuously changing meshes.
Model RMSE-1 (×$\times 10^{-2}$) RMSE-All (×$\times 10^{-2}$)
MGN (2021) 0.0618 24.08
EvoMesh 0.0544 7.412
Promotion 12.0% 69.2%
@@
We evaluate EvoMesh in a more challenging setting with time-varying meshes for paper folding simulation, generated using the ARCSim solver (Narain et al., 2012; Wu et al., 2023), and compare EvoMesh with MGN (Pfaff et al., 2021). Lino et al., BSMS-GNN, Eagle, and HCMT rely on pre-computed hierarchies during preprocessing, which limits their applicability in scenarios with dynamically changing mesh topologies. Therefore, we do not include them in this evaluation. We assess the models using ground-truth remeshing nodes provided by the ARCSim Adaptive Remeshing component, following the setup from (Pfaff et al., 2021). As shown in Table 3, EvoMesh achieves superior short-term and long-term accuracy compared to MGN, indicating that the time-evolving graph hierarchies in our approach can better fit physical systems with significant geometric variations, as represented by the time-varying input mesh structures.
@@
Due to the stochasticity of Gumbel-Softmax sampling in DiffSELECT, we evaluate the stability of trained EvoMesh by conducting three independent runs on the test set. The mean and standard deviations of the prediction errors reveal minimal discrepancies across different runs, as shown in Table 11 in Appendix C.6. These findings demonstrate that once trained, EvoMesh generates consistent graph hierarchies based on the same inputs.
@@ 4.3 Ablation Studies
EvoMesh has three key components: (i) evolving graph hierarchy, (ii) anisotropic intra-level propagation, (iii) learnable inter-level propagation. To evaluate the contribution of each component, we implement several ablated variants of EvoMesh, including: Static(Bi-stride)-Anisotropic-Unlearnable (M1), Static(Bi-stride)-Anisotropic-Learnable (M2), Uniform-Anisotropic-Learnable (M3), and Dynamic-Anisotropic-Unlearnable (M4), and compare them against the BSMS-GNN baseline, which uses static hierarchies, isotropic intra-level summation, and unlearnable inter-level propagation. Both M1 and M2 adopt the same static bi-stride hierarchy via preprocessing as BSMS-GNN. M3 constructs the hierarchy via uniform node sampling in each hierarchy, while M4 applies dynamic hierarchy construction without learnable inter-level updates.
@@
Figure 4: Ablation studies. We provide analyses of time-evolving hierarchies, anisotropic intra-level propagation, and learnable inter-level feature propagation. The red dashed lines represent results from BSMS-GNN (Cao et al., 2023). Lower values indicate better performance.
Figure 5: A demonstration of how the predicted anisotropic edge weights respond to dramatic changes in physical quantities over time. Top: Visualizations of the variance in the generated anisotropic weights, calculated on adjacent edges. Bottom: Variance in physical quantities over time. The strong correlation between them highlights the AMP’s ability to detect significant patterns in data.
Figure 4 demonstrates the effectiveness of direction-aware message propagation at both intra- and inter-levels, as well as the benefit of learning dynamic graph hierarchies. Comparing M1 with BSMS-GNN shows that integrating AMP improves performance even under a static hierarchy. Furthermore, the comparison between EvoMesh and M4, as well as between M1 with M2, highlights the importance of learnable inter-level propagation in capturing hierarchical signal flow. In addition, although M3 benefits from learnable propagation, its use of uniform node sampling results in suboptimal performance, suggesting the necessity of adaptive and data-aware hierarchy construction.
Furthermore, in Figure 5, we visualize the variance of predicted anisotropic edge weights and compare it with areas where physical quantities present substantial variations over time. The results reveal a strong correlation between the anisotropic learning mechanism and the rapidly changing dynamics of the physical system.
Table 4: One-step (RMSE-1) and 50-step (RMSE-50) rollout prediction errors on out-of-distribution (OOD) mesh resolutions. The average number of nodes in the test data is twice that of the training data, and the number of edges is three times greater.
Model RMSE-1 (×$\times 10^{-2}$) RMSE-50 (×$\times 10^{-2}$)
@@
MGN (2021) 1.0596 169.6 0.4215 0.0359 7.833 1829 55.96 0.2467
Lino et al. (2022) 25.893 144.4 0.8906 0.0475 65.21 1391 93.68 3.9845
BSMS-GNN (2023) 0.9177 202.3 0.6486 0.0474 2.097 1677 59.18 0.2554
HCMT (2024) 1.3864 205.5 1.0634 0.0354 7.541 2569 86.87 0.2957
EvoMesh 0.4855 126.7 0.5536 0.0368 1.077 812.5 58.29 0.3780
Table 5: Generalization results across domains with various scales of input velocities. The domain gap is presented by the variance and norm of data in training/test splits. Increase denotes the relative increase of the test data compared to the training data.
Split Cylinder Airfoil
Var Norm Var Norm
@@
Model Cylinder Airfoil
RMSE-1 RMSE-All RMSE-1 RMSE-All
MGN 4.99×$\times 10^{-3}$ 1.020 1.193 88.23
Lino et al. 5.60×$\times 10^{-3}$ 1.415 3.226 410.5
@@
Eagle 2.49×$\times 10^{-3}$ 0.273 0.931 51.83
HCMT 7.35×$\times 10^{-3}$ 1.047 1.697 63.18
EvoMesh 2.14×$\times 10^{-3}$ 0.091 0.665 22.57
Promotion 14.1% 63.7% 28.6% 25.6%
4.4 Generalization Analyses
Generalization to out-of-distribution mesh resolutions.
Nearly all existing machine learning models for mesh-based simulations are not resolution-free and may fail when evaluated on unseen mesh resolutions. We assess the generalization performance of EvoMesh by training it on low-resolution meshes and testing it on high-resolution meshes. The average number of nodes in the test data is twice that of the training data, and the number of edges is three times greater. Table 4 reports one-step and 50-step rollout errors on out-of-distribution (OOD) mesh resolutions. EvoMesh shows strong generalization on the CylinderFlow and Airfoil datasets, highlighting its zero-shot capability to handle refined mesh structures. This performance gain is largely attributed to EvoMesh’s ability to construct hierarchical graphs adaptively, allowing it to scale effectively with increased resolution. On the other hand, MGN achieves lower errors on the FlyingFlag and DeformingPlate datasets, indicating that mesh-based architectures remain effective for systems with more regular structures and smoother deformation dynamics. While our method does not yet achieve full generalization across arbitrary resolutions, truly resolution-free modeling remains an open challenge that calls for more advanced architectural design. Nevertheless, this holds significant value in practical applications and has the potential to greatly reduce the time overhead of numerical simulation processes for preparing the large-scale mesh data required for model training.
Generalization to physical variations.
We evaluate EvoMesh under strong distribution shifts in the input physical quantities. Table 5 presents data statistics and the RMSE results on the CylinderFlow and Airfoil datasets. EvoMesh consistently outperforms the compared models in both short-term and long-term simulations. This advantage mainly comes from its ability to learn evolving hierarchies and model intra-level and inter-level interactions based on physical context. When the fluid dynamics in the test set become more complex—characterized by increased variance in the velocity field over time—the dynamics patterns propagate more rapidly in space. EvoMesh adaptively constructs graph hierarchies and more effectively captures long-range node interactions in response to evolving physical contexts.